# On the relevance of pre-neural approaches in natural language processing pedagogy

### v1

@@ 6.2 Course Project
The group project is a centerpiece of the NLP-UNSW course, allowing students to explore diverse topics within NLP. Teams of four to five students are encouraged to select a specific area of interest, ranging from NLP topics of active research such as prompt recovery in large language models to purely applied topics like developing a virtual learning assistant using Retrieval Augmented Generation Lewis et al. (2020). To guide students effectively, we provide a detailed scope guideline document. The document outlines the following key aspects alongside their respective credits (cr):
@@
Findings:
The diversity of modeling algorithms is particularly interesting across the submissions (see Fig. 5). We note that pre-neural techniques such as such as lexicon-based systems, traditional machine learning algorithms (e.g., SVM, Naive Bayes), and feature engineering-based techniques serve towards foundation building, interpretable analyses and resource efficiency all while allowing the students to appreciate the advancements brought about by deep learning in NLP by benchmarking against these. Deep learning models (Fig. 5(a)), on the other hand, help achieve complex pattern recognition, state-of-art performance and end-to-end learning while helping the students understand the laws of scalability on large-scale datasets.
@@
(b) Pre-neural
Figure 5: Word clouds showing the key neural and pre-neural methods explored in the group projects.
In NLP-IITB, we list a few topics for students to choose from, while being open to new ideas. Providing topics helps to streamline project ideas while giving student teams a headstart. The projects are evaluated using a combination of demonstration and presentation, including an in-person discussion with the course team. The focus is the selection of the ‘right’ algorithm. The difference in the level of detail appears to reflect the experience of the two academics running the courses.
@@ 6.3 Tutorials
Tutorials are a weekly activity in NLP-UNSW (while not a part of NLP-IITB). In a typical tutorial, we first cover a focused review of the content covered in the lectures. We initially provide a brief overview of pre-neural methods, as these approaches often greatly assist students in understanding the task. Among the pre-neural topics, students show interest in learning about POS tagging and HMM-related algorithms. In the tutorial, we emphasize the mathematical derivation of HMM itself to improve the student’s understanding. Subsequently, we extend the discussion by integrating the latest research related to the lecture topics. This part is focused on expanding on the lecture content to cover recent, trending papers that are not included in the syllabus. For instance, when discussing Parameter-Efficient Fine-Tuning (PEFT), we incorporate a paper-reading session of the paper by He et al. (2021), explaining its essence from a unified perspective and how it can be generally summarized. This allows us to meet the expectations of students who may be interested in advanced topics. Next, we utilize code demonstrations to illustrate the concepts presented in the lectures, facilitating a deeper understanding for the students. Following this, we explain and demonstrate the coursework requirements, such as homework assignments. Finally, we conclude by addressing student questions, which may pertain to projects or other assignments. Particularly in the tutorials, the students show less interest in the pre-neural sections, focusing more on deep learning-based methods that are perceived as more employable. Students who actively participate and show interest in the derivation of various model principles are often also interested in research.
7 Making the case for pre-neural approaches
@@
Based on the considerations discussed so far, we find that the following arguments make the case for pre-neural approaches. Designing a curriculum based on course expectations along these arguments may help determine the ‘balance’ that is the focus of this paper.
Figure 6: Intuition-building using pre-neural approaches.
  1. 1.
@@
    Intuition-building: Rule-based and statistical approaches are great for intuition building and also to appreciate that NLP is challenging. Classroom exercises where students describe rules for a particular NLP task enable them to see why a good rule-based system would be laborious due to the well-known high-precision-low-recall setup. Statistical approaches help to highlight the importance of probability in NLP. Probability is the cornerstone of NLP models, including modern models. Softmax being the center of Transformer-based models is an example of that. Starting with Bayes theorem, a teacher can effectively tease out dependencies between different variables, serving to be a great explanation to lead into attention and related concepts. A pre-course survey done in NLP-IITB revealed that, out of 150 students, 80% preferred that the course build intuition as opposed to provide information. In fact, both the courses combine the pillars illustrated in  6. Linguistics motivates the problem, probability and code run as common threads allowing a comparison between the two kinds of approaches.
  2. 2.
@@
    Student Motivation: Several blog articles on NLP are available online. A clear differentiator in a university-based course is providing an inspiration to students. Towards this, pre-neural approaches are effective to inspire students to take up NLP projects. The challenges of rule-based systems are clear with some examples. Covering pre-neural approaches serves as a good motivator students to see why deep learning approaches have been revolutionary to the field. This feeling of inspiration is often reported in NLPb.
  3. 3.
@@
    Popular classical approaches: Pre-neural approaches have done well for linguistic tasks, particularly at the lower levels of the NLP hierarchy. Approaches like HMM and CRF work well for POS tagging and other token classification tasks. We observe that past courses
  4. 4.
@@
    Annotation: Pre-neural techniques, particularly in terms of annotation are the benchmark for modern NLP models. Tag sets were defined decades ago, but are still useful for their linguistic rigour. An example is the POS tagset. Nuances in Penn TreeBank about tags for "JJ" versus "JJR" (adjectives and comparative adjectives) help the students understand why they were designed in a certain way. This builds a linguistic mindset. The pre-course survey in NLPb revealed that 67% students preferred fundamental approaches versus state-of-the-art. The course covered a combination of the two.
  5. 5.
@@
    Cognitive load theory: Pre-neural techniques often involve explicit, well-defined rules and simpler models that can be more transparent and interpretable than their neural counterparts. This clarity can reduce extraneous cognitive load for learners by providing simpler examples of how inputs are transformed into outputs. For instance, a decision tree for language processing allows learners to see the exact paths through which decisions are made, thus aligning with the segmenting principle of cognitive load theory.
@@ 8 Conclusion
This paper investigated the role of pre-neural approaches in an introductory NLP course, by comparing two courses: one which was offered for the first time and another which has been running for ninteen years while being continually revised. We discussed key considerations and reasons why pre-neural approaches may be helpful, in terms of the past textbooks, lectures, and assessments. An understanding of pre-neural NLP not only builds a strong foundation for the students, but also enables them to look back at some of these methods to develop new techniques. We hope that the analyses presented in the paper will allow NLP educators and students alike to find the right balance between the pre-neural and deep learning approaches.

### v2

@@ 6.2 Course Project
Running in its first year, NLP-UNSW provides detailed guidelines for the course project. In contrast, NLP-IITB provides less guidelines but involves periodic consultation with the course team.
NLP-UNSW: The group project is a centerpiece of the NLP-UNSW course, allowing students to explore diverse topics within NLP. Teams of four to five students are encouraged to select a specific area of interest, ranging from NLP topics of active research such as prompt recovery in large language models to purely applied topics like developing a virtual learning assistant using Retrieval Augmented Generation Lewis et al. (2020). To guide students effectively, we provide a detailed scope guideline document. The document outlines the following key aspects alongside their respective credits (cr):
@@
Findings.
The diversity of modeling algorithms is particularly interesting across the submissions (see Fig. 5). We note that classical techniques such as such as lexicon-based systems, traditional machine learning algorithms (e.g., SVM, Naive Bayes), and feature engineering-based techniques serve towards foundation building, interpretable analyses, and resource efficiency all while allowing the students to appreciate the advancements brought about by deep learning in NLP by benchmarking against them. Deep learning models (Fig. 5(a)), on the other hand, help achieve complex pattern recognition, state-of-art performance, and end-to-end learning while helping the students understand the laws of scalability on large-scale datasets.
@@
(b) classical
Figure 5: Word clouds showing the key deep learning and classical methods explored in the group projects.
NLP-IITB: In NLP-IITB, we list a few topics for students to choose from, while being open to new ideas. Providing topics helps to streamline project ideas while giving student teams a headstart. The projects are evaluated using a combination of demonstration and presentation, including an in-person discussion with the course team. The focus is the selection of the ‘right’ algorithm. The difference in the level of detail appears to reflect the experience of the two academics running the courses.
@@ 6.3 Tutorials
Tutorials are a weekly activity in NLP-UNSW (while not a part of NLP-IITB). In a typical tutorial, we first cover a focused review of the content covered in the lectures. We initially provide a brief overview of classical methods, as these approaches often greatly assist students in understanding the task. Among the classical topics, students show interest in learning about POS tagging and HMM-related algorithms. In the tutorial, we emphasize the mathematical derivation of HMM itself to improve the student’s understanding. Subsequently, we extend the discussion by integrating the latest research related to the lecture topics. This part is focused on expanding on the lecture content to cover recent, trending papers that are not included in the syllabus. For instance, when discussing Parameter-Efficient Fine-Tuning (PEFT), we incorporate a paper-reading session of the paper by He et al. (2021), explaining its essence from a unified perspective and how it can be generally summarized. This allows us to meet the expectations of students who may be interested in advanced topics. Next, we utilize code demonstrations to illustrate the concepts presented in the lectures, facilitating a deeper understanding for the students. Following this, we explain and demonstrate the coursework requirements, such as homework assignments. Finally, we conclude by addressing student questions, which may pertain to projects or other assignments. Particularly in the tutorials, the students show less interest in the classical sections, focusing more on deep learning-based methods that are perceived as more employable. Students who actively participate and show interest in the derivation of various model principles are often also interested in research.
7 Making the Case for Classical Approaches
@@
Based on the considerations discussed so far, we are now able to make several arguments that may influence the inclusion of classical approaches in NLP courses. Designing a curriculum based on course expectations along these arguments may help determine the ‘balance’ that is the focus of this paper.
Figure 6: Intuition-building using classical approaches.
7.1 Intuition-building
@@
The conveners of NLP-UNSW and NLP-IITB find that rule-based and statistical approaches are great for intuition building and also appreciate that NLP is challenging. Classroom exercises where students describe rules for a particular NLP task enable them to see why a good rule-based system would be laborious due to the well-known high-precision-low-recall setup. Statistical approaches help to highlight the importance of probability in NLP. Probability is the cornerstone of NLP models, including modern models. Softmax being the center of Transformer-based models is an example of that. Starting with Bayes theorem, a teacher can effectively tease out dependencies between different variables, serving to be a great explanation to lead to attention and related concepts, especially during lectures.
A pre-course survey done in NLP-IITB revealed that, out of 150 students, 80% preferred that the course build intuition as opposed to providing information. Information refers to providing a suite of approaches and methods, while intuition refers to the motivation (linguistic or mathematical, etc.) underlying the methods. In fact, both courses combine the pillars illustrated in  6. Linguistics motivates the problem, probability and code run as common threads allowing a comparison between the two kinds of approaches. This motivation is highlighted by the instructor of NLP-IITB, by leveraging examples from his rich research experience. In contrast, the instructor of NLP-UNSW prefers to give examples from industry applications, given their professional experience in the tech industry.
7.2 Student Motivation
@@
Several blog articles on NLP are available online. A clear differentiator in a university-based course is inspiring students using blended learning Deng and Yuen (2010). Towards this, classical approaches are effective in inspiring students to take up NLP projects. The conveners of NLP-UNSW and NLP-IITB observe that the challenges of rule-based systems are clear with some examples, during lectures. Covering classical approaches serves as a good motivator for students to see why deep learning approaches have been revolutionary to the field. This feeling of inspiration is often reported in NLP-IITB.
7.3 Popular classical approaches
@@
Classical approaches have done well for linguistic tasks, particularly at the lower levels of the NLP hierarchy. Approaches like HMM and CRF work well for POS tagging and other token classification tasks. We observe that past courses also tend to cover classical methods as an introduction to deep learning methods. Comparison of popular classical approaches with deep learning approaches can aid learning during projects and assignments, as described in Section 6.
7.4 Annotation
@@
Classical techniques, particularly in terms of annotation, are the benchmark for modern NLP models since tag sets used in classical approaches are still useful for their linguistic rigour. An example is the POS tagset. Nuances in Penn TreeBank about tags for “JJ" versus “JJR" (adjectives and comparative adjectives) help the students understand why they were designed in a certain way. The pre-course survey in NLP-IITB revealed that 67% students preferred fundamental approaches versus state-of-the-art. The course covered a combination of the two.
7.5 Cognitive load theory
@@
Classical techniques often involve explicit, well-defined rules and simpler models that can be more transparent and interpretable than their neural counterparts. This clarity can reduce extraneous cognitive load for learners by providing simpler examples of how inputs are transformed into outputs. For instance, a decision tree for language processing allows learners to see the exact paths through which decisions are made, thus aligning with the segmenting principle of cognitive load theory.
@@ 8 Conclusion
Classical approaches for NLP refer to those prior to the predominant use of neural networks and deep learning. This paper investigated the role of classical approaches in an introductory NLP course, by comparing perspectives from two courses: one which was offered for the first time and another which has been running for nineteen years while being continually revised. We discussed key considerations and reasons why classical approaches may be helpful, in terms of the past textbooks, lectures, and assessments. An understanding of classical NLP not only builds a strong foundation for the students but also enables them to look back at some of these methods to develop new techniques. We hope that the analyses presented in the paper will allow NLP educators and students alike to find the right balance between the classical and deep learning approaches.
Limitations
The books and courses described in the paper are a subset curated only as an example. It does not represent a complete list. Similarly, the sections derive from insights teaching the two courses covered in the paper, and are not necessarily prescriptive.
Ethical Considerations
The paper describes summary statistics of student surveys without identifying individual students or student cohorts. The assessment section provides only a high-level view of the assignments and lectures, and may not result in implications to academic integrity in the future versions of the courses.
Acknowledgment
The authors thank Dipankar Srirag, UNSW Sydney, for his feedback on a near-camera-ready draft of the paper.